OpenStudy (jessiegonzales):

MEDAL!!!! Name the property of real numbers illustrated by the equation. -2(x+11)=-2x-22 A. Associative Property of Multiplication B. Distributive Property C. Commutative Property of Addition D. Associative Property of Addition

OpenStudy (jessiegonzales):

@zepdrix

OpenStudy (owlcoffee):

Which one do you think it is?

OpenStudy (jessiegonzales):

C

OpenStudy (owlcoffee):

No, the commutative property, states that if any number added to an other it'll be the same if we turn it around: \[a+b=b+a\]

OpenStudy (jessiegonzales):

OH...... A

OpenStudy (owlcoffee):

Wrong again, the associative property states that the group of two number aded to another, is the same as any integer of the group, added to any of the numbers: \[a+(b+c)=(a+b)+c\]

OpenStudy (jessiegonzales):

sORRY final answer B

OpenStudy (owlcoffee):

Correct!

OpenStudy (jessiegonzales):

THANKYOU! :)

OpenStudy (jessiegonzales):

Mind helping with another?

OpenStudy (owlcoffee):

sure, make an other post and I'll be there as soon as I can.